OpenStudy (anonymous):

Solve: −5(x − 2) = 15

OpenStudy (anonymous):

-5x + 10 = 15 -5x = 5 x = -1

OpenStudy (pokemon23):

dam you panda :(

OpenStudy (anonymous):

haha

OpenStudy (pokemon23):

grrr

OpenStudy (anonymous):

I'm too fast for you?

OpenStudy (anonymous):

I give you medal anyway (:

OpenStudy (anonymous):

you got 2 medals and I got only 1 ok?

OpenStudy (anonymous):

Lmao I didn't know you could give out more than one xD